A bag contains 4 red and 3 blue balls. John and Paul selects a ball at random and replaces it. The experiment is repeated 7 times. Find the probability that:
a) all 7 selections are red
b) at least 5 are red
c) no more than 2 are red

Answer by edjones(8007)

This is a binomial distribution because there is replacement.
x=4/7, y=3/7
.
a) (4/7)^7=.020
.
b) x^7+7x^6y+21x^5y^2=.359
.
c) 21x^2y^5+7xy^6+y^7=.127
.
Ed